OpenStudy (anonymous):

A beam in a viewing booth for officials at the Daytona 500 auto race is being replaced. The strength(weight) of a beam (s) is directly proportional to its width (w) and the square of its depth (d). Write an equation for s involving the width and the depth that could be used to solve for the constant of variation k. s=k*w*d^2 The current beam is 3.5 inches wide and 6 inches deep and will support a weight of 1260 pounds. Find the constant of variation k. 1260=k*3.5*6^2 1260=k*126 k=1260/126 k=10 If a replacement beam is 3 inches wide and 7 inches deep, how much weight can it suppo

OpenStudy (anonymous):

Aby what is the trouble here.You have found out everything.you know k,the equation relating depth ,length and weight.Substitute corresponding values.i.e here find s. Yopu know s=10*w*d^2 So weight supported =s=10*3*7^2

OpenStudy (anonymous):

I have been working on different math assignments for the past 4 hours, just wanted to be sure I was answering this correctly. Thank you for your help and reassurance.